OpenStudy (mallorysipp234):

why arent x-4 and x-4 inverses of each other???

OpenStudy (anonymous):

Assuming you mean "inverse" as in the inverse of a function, suppose \(f(x)=x-4\) is indeed the same as its inverse, \(f^{-1}(x)\). Then it must have the property that \(f\left(f^{-1}(x)\right)=f^{-1}\left(f(x)\right)=x\). But this clearly isn't the case, since \[f(x-4)=(x-4)-4=x-8\neq x\]

jhonyy9 (jhonyy9):

@mia99 the last term not is +8 bc (-4)*(-4)=16

jhonyy9 (jhonyy9):

and so (x-4)*(x+4) not equal x^2 -8 this is equal x^2 -16 ATTENTION please !!!

Directrix (directrix):

> why arent x-4 and x-4 inverses of each other??? I interpreted the reference to "inverse" as being additive inverse. A real number and its additive inverse sum to zero. x-4 and x-4 sum to 2x - 8 which is not zero. Therefore, they are not additive inverses.

Directrix (directrix):

>> why arent x-4 and x-4 inverses of each other??? Then, I thought that the question may refer to multiplicative inverses. A real number (excluding 0) and its multiplicative inverse have a product of one. (x - 4) * (x - 4) does not equal 1. (x - 4) and (x -4) are not multiplicative inverses.